If the width x and height y of a rectangle are both changed in a way that the area remains constant at sixteen units.
Determine the perimeter P of the rectangle as a function of the width.

Answer by TimothyLamb(4379)

a = Lw = 16
L = 16/w
---
p = 2L + 2w
p = 2(16/w) + 2w
---
Answer:
p = 32/w + 2w
where w cannot be zero
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php